OpenStudy (anonymous):

TRUE OR FALSE. The slope of the line that passes through (-2, 6) and (1, 15) is 3.

Directrix (directrix):

True

OpenStudy (anonymous):

True...because the slipope is tha same

OpenStudy (anonymous):

true :)

OpenStudy (anonymous):

m = ( 15 -6) / ( 1 + 2) = 9 /3 = 3